# Barycentric Neural Networks and Length-Weighted Persistent Entropy Loss: A Green Geometric and Topological Framework for Function Approximation

### v2

@@ 3 Lenght-Weighted Persistent Entropy
Computing persistent entropy using lower star filtration, is suitable for comparing discrete piecewise linear functions and analyzing time series data [25]. However, standard persistent entropy presents a fundamental limitation: two different functions can still share a similar persistent entropy, as the entropy only encodes a global distribution of bar lengths in the persistence diagram, not taking into account its length. To address this issue, we propose a new persistent entropy variant: the length-weighted persistent entropy ($\operatorname{LWPE}$). Formally, we define $\operatorname{LWPE}$ as:
@@
This formulation maintains the entropy structure but introduces a crucial weighting mechanism based on the length of the bars. While $\operatorname{PE}$ captures the uncertainty in bar distributions, $\operatorname{LWPE}$ incorporates the total bar lengths, giving more weight to longer, topologically significant features.
Figure 2: Comparison of persistent entropy (PE) and length-weighted persistent entropy (LWPE) on three subsampled versions of the sine function. The corresponding PE and LWPE values are reported in the title of each subplot.
Figure 2 illustrates standard persistent entropy limitation with three different subsampled versions of the same sine function. Although the point-cloud representations exhibit clear differences in resolution, the values of PE remain remarkably close. This occurs because PE primarily captures the global balance of bar lengths rather than their absolute scale. For instance, in the middle plot, the function is represented with very few points, but since the relative proportions of the peaks are preserved, the persistence diagram remains proportionally similar to that of the original function, resulting in a nearly identical PE value despite the drastic reduction in resolution. In contrast, as shown in the same plot, LWPE reveals that such a coarse representation is insufficient: it is not enough to preserve proportional bar lengths in the diagram, but rather a similar peak structure is also required to achieve comparable values. This is why the right plot, where both the proportions and the peaks are consistent with the original, yields LWPE values that are closer, while the middle plot does not. As we will see later in Section 5, using LWPE in our framework leads to significantly better approximation quality in practice than using classical persistent entropy.
4 Barycentric Neural Networks (BNNs)
Generally speaking, a neural network can be viewed as a parametrized mathematical function $N_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{m}$, where $\theta$ represents the set of parameters that define the network architecture. In an abuse of notation, such set of parameters includes weights, bias, activation functions, and other architectural considerations like the number and type of layers. A general introduction to neural networks can be found in [13].
In this paper, we focus on one of the simplest types of artificial neural networks, the multi-layer feed-forward networks, which consist of sequences of stacked densely connected layers including input, hidden and output layers. The operation of a hidden layer can be mathematically expressed as: $y=h(Wx+b)$, where $x\in\mathbb{R}^{d}$ is the input vector, $W\in\mathbb{R}^{m\times d}$ is the weight matrix, $b\in\mathbb{R}^{m}$ is the bias vector, and $h$ is a nonlinear activation function (e.g., ReLU).
@@ 4 Barycentric Neural Networks (BNNs)
$\operatorname{BNN}_{\sigma}\big{(}(t_{0},t_{1},\dots,t_{d})\big{)}=\sum_{i=0}^{d}\operatorname{ReLU}\Big{(}1-\operatorname{ReLU}(1-t_{i})-\sum_{j=0}^{d}\big{(}\operatorname{step}^{*}(-t_{j})+\operatorname{step}^{*}(t_{j}-1)\big{)}\Big{)}\cdot g(v_{i})\,$,
where $(t_{0},t_{1},\dots,t_{d})$ are the barycentric coordinates of a given $p\in\mathbb{R}^{d}$ with respect to $\sigma$, ReLU corresponds to the activation function defined by $\operatorname{ReLU}(t)=\max\{0,t\}$, for $t\in\mathbb{R}$, and $\operatorname{step}^{*}$ is the activation function defined as follows:
$\operatorname{step}^{*}(t)=1-\operatorname{step}(-t)=\left\{\begin{array}[]{cl}1&\mbox{ if $t>0$}\\ 0&\mbox{ if $t\leq 0$,}\end{array}\right$.
where $\operatorname{step}(t)=1$ if $t\geq 0$ and 0 otherwise.
@@
    If $p\in|\sigma|$ then $\operatorname{BNN}_{\sigma}\big{(}(t_{0},t_{1},\dots,t_{d})\big{)}=\sum_{i=0}^{d}t_{i}\cdot g(v_{i})$.
@@
    If $p\not\in|\sigma|$ then $\operatorname{BNN}_{\sigma}\big{(}(t_{0},t_{1},\dots,t_{d})\big{)}=0$.
@@
    We then have that $\operatorname{BNN}_{\sigma}\big{(}(t_{0},t_{1},\dots,t_{d})\big{)}=\sum_{i=0}^{d}t_{i}\cdot g(v_{i})$.
  • (b)
    If $p\not\in|\sigma|$ then, by the definition, the barycentric coordinates $(t_{0},t_{1},\dots,t_{d})$ of $p$ with respect to $\sigma$ satisfy that there is a non-emtpy set $J\subset\{0,1,\dots,d\}$ such that for any $j\in J$, $t_{j}<0$ or $t_{j}>1$. In that case, $\operatorname{step}^{*}(-t_{j})=1$ and $\operatorname{step}^{*}(1-t_{j})=0$, or $\operatorname{step}^{*}(-t_{j})=0$ and $\operatorname{step}^{*}(t_{j}-1)=1$. Therefore, $\operatorname{step}^{*}(-t_{j})+\operatorname{step}^{*}(t_{j}-1)=1$. Then,
@@
    for all $i\in\{0,1,\dots,d\}$, concluding that $\operatorname{BNN}_{\sigma}\big{(}(t_{0},t_{1},\dots,t_{d})\big{)}=0$.
@@
$\operatorname{BNN}_{\sigma}\big{(}(1-t,t)\big{)}$ $=\operatorname{ReLU}\big{(}1-\operatorname{ReLU}(t)-(\operatorname{step}^{*}(t-1)+\operatorname{step}^{*}(-t)\ +\operatorname{step}^{*}(-t)+\operatorname{step}^{*}(t-1)\;)\big{)}\cdot g(a)$
@@
Therefore, $\operatorname{BNN}_{\sigma}\big{(}(1-t,t)\big{)}=0$ if $t<0$ or $t>1$, that is, if $x$ is outside the closed interval $[a,b]$, ensuring locality.
Figure 3: Structure and parameters of $\operatorname{BNN}_{\sigma}$ in the case of $\sigma=\{a,b\}$, with $a<b\in\mathbb{R}$. Each hidden neuron is denoted as $h_{ij}$, where $i$ indicates the hidden layer and $j$ the neuron index within that layer (e.g., $h_{22}$ is the second neuron of the second hidden layer). The numbers written above the neurons correspond to biases, while the labels below indicate the activation functions. The numbers on the edges represent weights.
As we can see in Fig. 3, in the case of $\sigma=\{a,b\}$ with $a<b\in\mathbb{R}$, $\operatorname{BNN}_{\sigma}$ corresponds to a multi-layer feedforward network with a single-neuron input and output layer (identity activation, referred to as $id$). It has two hidden layers with four and two neurons, respectively, where neurons in the same layer may have different activation functions.
@@
$\mathrm{BNN}(p)=\frac{1}{\#\{k:\mathrm{BNN}_{\sigma_{k}}\big{(}(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})\big{)}>0\}}\cdot\sum_{k=0}^{n}\mathrm{BNN}_{\sigma_{k}}\big{(}(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})\big{)},$
where $\{k:\mathrm{BNN}_{\sigma_{k}}\big{(}(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})\big{)}>0\}$ counts the number of local networks $\mathrm{BNN}_{\sigma_{k}}$ that are nonzero at the barycentric coordinates $(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})$ of $p$ with respect to $\sigma^{k}$.
In other words, the value of $\mathrm{BNN}$ at a point $p$ is the average of the values of the local networks $\mathrm{BNN}_{\sigma_{k}}$ for which $p$ lies in $|\sigma_{k}|$, which is precisely detected by $\mathrm{BNN}_{\sigma_{k}}\big{(}(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})\big{)}>0$. This normalization properly accounts for overlapping simplices: points at the intersection of multiple simplices (e.g., shared vertices or edges) yield multiple nonzero local evaluations, and the $\operatorname{BNN}$ formula averages them to ensure continuity and consistency.
@@
$\operatorname{BNN}_{\sigma_{1}}\big{(}(0,1)\big{)}=$ $\operatorname{ReLU}\big{(}1-\operatorname{ReLU}(1)-2\cdot\operatorname{step}^{*}(0)-2\cdot\operatorname{step}^{*}(-1)\big{)}\cdot g(a_{1})$
@@
$\operatorname{BNN}_{\sigma_{2}}\big{(}(1,0)\big{)}=$ $\operatorname{ReLU}\big{(}1-\operatorname{ReLU}(0)-2\cdot\operatorname{step}^{*}(-1)-2\cdot\operatorname{step}^{*}(0)\big{)}\cdot g(b_{1})$
@@
Then, $\operatorname{BNN}_{\sigma_{1}}\big{(}(0,1)\big{)}+\operatorname{BNN}_{\sigma_{2}}\big{(}(1,0)\big{)}=g(a_{2})+g(b_{1})=2\cdot g(b_{1})=2\cdot g(a_{2})$ (since $b_{1}=a_{2})$. This illustrates the redundancy at joining points and motivates the normalization in the global BNN.
@@
    $\operatorname{BNN}(p)=\operatorname{BNN}_{\sigma^{k}}\big{(}(t_{0}^{k},t_{1}^{k},\dots,t_{d}^{k})\big{)}=\sum_{i=0}^{d}t^{k}_{i}\cdot g(v^{k}_{i}).$
@@
First, recall the definition of $\operatorname{BNN}_{\sigma}$ for a given point $p\in\mathbb{R}^{d}$ with barycentric coordinates $(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})$ with respect to a simplex $\sigma^{k}$:
$\operatorname{BNN}_{\sigma_{k}}\big{(}(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})\big{)}=\sum_{i=0}^{d}\operatorname{ReLU}\Big{(}1-\operatorname{ReLU}(1-t^{k}_{i})-\sum_{j=0}^{d}\big{(}\operatorname{step}^{*}(-t^{k}_{j})+\operatorname{step}^{*}(t^{k}_{j}-1)\big{)}\Big{)}\cdot g(v^{k}_{i}).$
Now, let us prove the statements.
  • (1)
    By definition, the barycentric coordinates of $v^{k}_{\ell}$ with respect to $\sigma_{k}$ are $(0,\stackrel{{\scriptstyle\ell}}{{\dots}},0,1,0\stackrel{{\scriptstyle d-\ell}}{{\dots}},0)$ where the value $1$ is at position $\ell$. Since $\operatorname{step}^{*}(0)=step^{*}(-1)=0$ and $1-\operatorname{ReLU}(1-0)=0$ then
    $\operatorname{BNN}_{\sigma_{k}}\big{(}(0,\stackrel{{\scriptstyle\ell}}{{\dots}},0,1,0\stackrel{{\scriptstyle d-\ell}}{{\dots}},0)\big{)}=\operatorname{ReLU}\Big{(}1-\operatorname{ReLU}(1-1)\Big{)}\cdot g(v^{k}_{\ell})=g(v^{k}_{\ell}).$
@@
    $BNN_{\sigma_{k^{\prime}}}\big{(}(0,\stackrel{{\scriptstyle\ell^{\prime}}}{{\dots}},0,1,0\stackrel{{\scriptstyle d-\ell^{\prime}}}{{\dots}},0)\big{)}=g(v^{k^{\prime}}_{\ell^{\prime}})=g(v^{k}_{\ell}).$
    We can then conclude that, $\operatorname{BNN}(v^{k}_{\ell})=g(v^{k}_{\ell}).$
  • (2)
@@
    $\operatorname{BNN}_{\sigma_{k}}\big{(}(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})\big{)}=\sum_{i=0}^{d}t^{k}_{i}\cdot g(v^{k}_{i}).$
    If $p\in|\sigma^{i}|\cap|\sigma^{j}|$ for some $i,j\in\{0,1,\dots,n\}$, let $(t^{i}_{0},t^{i}_{1},\dots,t^{i}_{d})$ and $(t^{j}_{0},t^{j}_{1},\dots,t^{j}_{d})$ be the barycentric coordinates of $p$ with respect to $\sigma_{i}$ and $\sigma_{j}$ respectively. By the definition of simplicial complexes, there is an $\ell$-simplex $\sigma=\{v_{0},v_{1},\dots,v_{\ell}\}\subset\sigma_{i},\sigma_{j}$ with $\ell<d$ such that $p\in|\sigma|$ and satisfying that there are sets of indices $I,J\subset\{0,1,\dots,d\}$ such that $v_{r}=v^{i}_{r_{i}}=v^{j}_{r_{j}}$ and $t_{r}=t^{i}_{r_{i}}=t^{j}_{r_{j}}$ for some $r_{i}\in I$ and $r_{j}\in J$ and for all $r\in\{0,1,\dots,\ell\}$, being $(t_{0},t_{1},\dots,t_{\ell})$ the barycentric coordinates of $p$ with respect to $\sigma$. Putting all together, we have that
    $\operatorname{BNN}_{\sigma^{i}}\big{(}(t_{0}^{i},t_{1}^{i},\dots,t_{d}^{i})\big{)}=\sum_{r_{i}\in I}t^{i}_{r_{i}}\cdot g(v^{i}_{r_{i}})=\sum_{r_{j}\in I}t^{j}_{r_{j}}\cdot g(v^{j}_{r_{j}})=BNN_{\sigma^{j}}\big{(}(t_{0}^{j},t_{1}^{j},\dots,t_{d}^{j})\big{)}.$
    We then conclude that, for any $k\in\{0,1,\dots,n\}$ with $p\in|\sigma_{k}|$,
    $\operatorname{BNN}(p)=\sum_{i=0}^{d}t^{k}_{i}\cdot g(v^{k}_{i})$
    where $(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})$ are the barycentric coordinates of $p$ with respect to $\sigma_{k}$.
  • (3)
    If $p\not\in C$ then, $p\not\in|\sigma_{k}|$ for any given $k\in\{0,1,\dots,n\}$. Therefore, by Lemma 1, $\operatorname{BNN}_{\sigma^{k}}\big{(}(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})\big{)}=0$ and so $\operatorname{BNN}(p)=0$.
  • (4)
@@
    where $m_{k}\in\mathbb{R}^{d}$ and $c_{k}\in\mathbb{R}$, concluding the proof.
@@
Any $\operatorname{CPLF}$ $g:[A,B]\to\mathbb{R}$ can be represented by a $\operatorname{BNN}$, with the key property that the $\operatorname{BNN}$ does not require an analytic definition. Instead, to specify the $\operatorname{CPLF}$ or the corresponding $\operatorname{BNN}$, it is sufficient to provide a finite set of transition (base) points $\{a_{i}\}\subset\mathbb{R}$, where the $\operatorname{CPLF}$ shifts between linear segments, together with the $\operatorname{CPLF}$ value at these points. This allows the $\operatorname{BNN}$ to capture the contribution of each segment relative to the input, thereby generating the correct output. For consistency with the terminology introduced in the Introduction, in the following we shall refer to these transition points as base points.

### v3

@@ 3 Lenght-Weighted Persistent Entropy
Persistent entropy ($\operatorname{PE}$) computed via lower-star filtration is suitable for comparing discrete piecewise linear functions and analyze time series [26]. However, PE suffers from a fundamental limitation: it is invariant under uniform rescaling of the function values. This occurs because $\operatorname{PE}$ depends only on the relative proportions of bar lengths in the persistence diagram, not their absolute magnitudes. Consequently, two functions with markedly different scales or resolutions can yield nearly identical $\operatorname{PE}$ values, even when their geometric or topological fidelity differs substantially.
To overcome this limitation,we introduce length-weighted persistent entropy ($\operatorname{LWPE}$), a variant that preserves the entropy structure while incorporating absolute persistence information. Formally, $\operatorname{LWPE}$ is defined as:
@@
Unlike $\operatorname{PE}$, which is scale-invariant, $\operatorname{LWPE}$ explicitly weights each term by the absolute persistence $\ell_{i}$. This ensures that longer-lived (i.e., more topologically significant) features contribute more substantially to the overall entropy, making $\operatorname{LWPE}$ sensitive to both the distribution and the scale of topological features.
Figure 2: Comparison of persistent entropy ($\operatorname{PE}$) and length-weighted persistent entropy ($\operatorname{LWPE}$) on three subsampled versions of a sine function. Despite similar $\operatorname{PE}$ values across resolutions, $\operatorname{LWPE}$ distinguishes coarse approximations (middle) from faithful ones (right), reflecting differences in absolute topological scale.
Figure 2 illustrates this difference. Three subsampled versions of the same sine function are shown. Although the relative shape of the signal is preserved, the middle plot uses so few points that the resulting piecewise linear approximation poorly captures the true peak structure. Standard $\operatorname{PE}$ remains nearly unchanged across all three cases because it only reflects the normalized distribution of bar lengths. In contrast, $\operatorname{LWPE}$ drops significantly in the middle case, correctly indicating a loss of topological fidelity due to reduced resolution. Only when both the relative proportions and the absolute scale of features are preserved (right plot) does $\operatorname{LWPE}$ remain close to the reference value.
As shown later in Section 5, using a topological loss function based on $\operatorname{LWPE}$ leads to substantially improved approximation quality, particularly in low-resource settings, by guiding the optimization of base points toward solutions that preserve not only the shape but also the scale of topologically relevant structures.
4 Barycentric Neural Networks
A neural network is typically viewed as a parametrized mathematical function $N_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{m}$, where $\theta$ encompasses weights, biases, activation functions, and architectural choices such as depth and layer types [13]. In this paper, we focus on multi-layer feed-forward networks, whose hidden layers compute transformations of the form $y=h(Wx+b)$, with input $x\in\mathbb{R}^{d}$, weight matrix $W$, bias vector $b$, and nonlinear activation $h$ (e.g., ReLU).
We now introduce the Barycentric Neural Network ($\operatorname{BNN}$), a shallow architecture that leverages barycentric coordinates to represent continuous piecewise linear functions exactly. The key idea is to replace trainable weights with geometrically meaningful base points, turning the learning problem into an optimization over point positions.
@@ 4 Barycentric Neural Networks
$\operatorname{BNN}_{\sigma}(p)=\sum_{i=0}^{d}\operatorname{ReLU}\Big(1-\operatorname{ReLU}(1-t_{i})-\sum_{j=0}^{d}\big(\operatorname{step}^{*}(-t_{j})+\operatorname{step}^{*}(t_{j}-1)\big)\Big)\cdot g(v_{i})\,$,
where $p\in\mathbb{R}^{d}$, $t_{i}=\frac{\lambda_{i}}{\sum_{j=0}^{d}\lambda_{j}}$, $\lambda_{i}=\det\left(\begin{array}[]{ccccccc}v_{0}&\cdots&v_{i-1}&p&v_{i+1}&\cdots&v_{d}\\ 1&\cdots&1&1&1&\cdots&1\end{array}\right)$, $i=0,1,\dots,d$, ReLU corresponds to the activation function defined by $\operatorname{ReLU}(t)=\max\{0,t\}$, for $t\in\mathbb{R}$, and $\operatorname{step}^{*}$ is the activation function defined as $\operatorname{step}^{*}(t)=1-\operatorname{step}(-t)=\left\{\begin{array}[]{cl}1&\mbox{ if $t>0$}\\ 0&\mbox{ if $t\leq 0$}\end{array}\right$. where $\operatorname{step}(t)=1$ if $t\geq 0$ and 0 otherwise.
@@
    If $p\in|\sigma|$ then $\operatorname{BNN}_{\sigma}(p)=\sum_{i=0}^{d}t_{i}\cdot g(v_{i})$.
@@
    If $p\not\in|\sigma|$ then $\operatorname{BNN}_{\sigma}(p)=0$.
@@
We then have that $\operatorname{BNN}_{\sigma}(p)=\sum_{i=0}^{d}t_{i}\cdot g(v_{i})$.
If $p\not\in|\sigma|$ then, by the definition, the barycentric coordinates $(t_{0},t_{1},\dots,t_{d})$ of $p$ with respect to $\sigma$ satisfy that there is a non-emtpy set $J\subset\{0,1,\dots,d\}$ such that for any $j\in J$, $t_{j}<0$ or $t_{j}>1$. In that case, $\operatorname{step}^{*}(-t_{j})=1$ and $\operatorname{step}^{*}(1-t_{j})=0$ when $t_{j}<0$, or $\operatorname{step}^{*}(-t_{j})=0$ and $\operatorname{step}^{*}(t_{j}-1)=1$ when $t_{j}>1$. Therefore, $\operatorname{step}^{*}(-t_{j})+\operatorname{step}^{*}(t_{j}-1)=1$. Then,
@@
for all $i\in\{0,1,\dots,d\}$, concluding that $\operatorname{BNN}_{\sigma}(p)=0$. ∎
@@
$\operatorname{BNN}_{\sigma}(x)$ $=\operatorname{ReLU}\big(1-\operatorname{ReLU}(t)-(\operatorname{step}^{*}(t-1)+\operatorname{step}^{*}(-t)\ +\operatorname{step}^{*}(-t)+\operatorname{step}^{*}(t-1))\big)\cdot g(a)$
@@
Therefore, $\operatorname{BNN}_{\sigma}(x)=0$ if $t<0$ or $t>1$, that is, if $x$ is outside the closed interval $[a,b]$, ensuring locality.
Figure 3: Structure and parameters of $\operatorname{BNN}_{\sigma}$ in the case of $\sigma=\{a,b\}$, with $a<b\in\mathbb{R}$, $x\in\mathbb{R}$. Each hidden neuron is denoted as $h_{ij}$, where $i$ indicates the hidden layer and $j$ the neuron index within that layer (e.g., $h_{22}$ is the second neuron of the second hidden layer). The numbers written above the neurons correspond to biases, while the labels below indicate the activation functions. The numbers on the edges represent weights. The first hidden layer only aims to compute the barycentric coordinate $t$ within the simplex $\sigma$ , while the subsequent layers determine the contribution of the point within that simplex.
As we can see in Fig. 3, in the case of $\sigma=\{a,b\}$ with $a<b\in\mathbb{R}$, $\operatorname{BNN}_{\sigma}$ corresponds to a multi-layer feedforward network with a single-neuron input and output layer (identity activation, referred to as $id$). It has three hidden layers with one, four and two neurons, respectively, where neurons in the same layer may have different activation functions.
@@
$\mathrm{BNN}(p)=\frac{1}{\#\{k:\mathrm{BNN}_{\sigma_{k}}(p)>0\}}\cdot\sum_{k=0}^{n}\mathrm{BNN}_{\sigma_{k}}(p),$
where $\{k:\mathrm{BNN}_{\sigma_{k}}(p)>0\}$ counts the number of local networks $\mathrm{BNN}_{\sigma_{k}}$ that are nonzero at the barycentric coordinates $(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})$ of $p$ with respect to $\sigma^{k}$.
In other words, the value of $\mathrm{BNN}$ at a point $p$ is the average of the values of the local networks $\mathrm{BNN}_{\sigma_{k}}$ for which $p$ lies in $|\sigma_{k}|$, which is precisely detected by $\mathrm{BNN}_{\sigma_{k}}(p)>0$. This normalization properly accounts for overlapping simplices: points at the intersection of multiple simplices (e.g., shared vertices or edges) yield multiple nonzero local evaluations, and the $\operatorname{BNN}$ formula averages them to ensure continuity and consistency.
@@
$\operatorname{BNN}_{\sigma_{1}}(a_{2})=$ $\operatorname{ReLU}\big(1-\operatorname{ReLU}(1)-2\cdot\operatorname{step}^{*}(0)-2\cdot\operatorname{step}^{*}(-1)\big)\cdot g(a_{1})$
@@
$\operatorname{BNN}_{\sigma_{2}}(b_{1})=$ $\operatorname{ReLU}\big(1-\operatorname{ReLU}(0)-2\cdot\operatorname{step}^{*}(-1)-2\cdot\operatorname{step}^{*}(0)\big)\cdot g(b_{1})$
@@
Then, $\operatorname{BNN}_{\sigma_{1}}(a_{2})+\operatorname{BNN}_{\sigma_{2}}(b_{1})=g(a_{2})+g(b_{1})=2\cdot g(b_{1})=2\cdot g(a_{2})$ (since $b_{1}=a_{2})$. This illustrates the redundancy at joining points and motivates the normalization in the global BNN.
@@
    $\operatorname{BNN}(p)=\operatorname{BNN}_{\sigma^{k}}(p)=\sum_{i=0}^{d}t^{k}_{i}\cdot g(v^{k}_{i}).$
@@
First, recall the definition of $\operatorname{BNN}_{\sigma}$ for a give point $p\in\mathbb{R}^{d}$ with barycentric coordinates $(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})$ with respect to a simplex $\sigma^{k}$:
$\operatorname{BNN}_{\sigma_{k}}(p)=\sum_{i=0}^{d}\operatorname{ReLU}\Big(1-\operatorname{ReLU}(1-t^{k}_{i})-\sum_{j=0}^{d}\big(\operatorname{step}^{*}(-t^{k}_{j})+\operatorname{step}^{*}(t^{k}_{j}-1)\big)\Big)\cdot g(v^{k}_{i}).$
Now, let us prove the statements. By definition, the barycentric coordinates of $v^{k}_{\ell}$ with respect to $\sigma_{k}$ are $(0,\stackrel{{\scriptstyle\ell}}{{\dots}},0,1,0\stackrel{{\scriptstyle d-\ell}}{{\dots}},0)$ where the value $1$ is at position $\ell$.Since $\operatorname{step}^{*}(0)=step^{*}(-1)=0$ and $1-\operatorname{ReLU}(1-0)=0$ then
$\operatorname{BNN}_{\sigma_{k}}(p)=\operatorname{ReLU}\Big(1-\operatorname{ReLU}(1-1)\Big)\cdot g(v^{k}_{\ell})=g(v^{k}_{\ell}).$
@@
$\operatorname{BNN}_{\sigma_{k^{\prime}}}(p)=g(v^{k^{\prime}}_{\ell^{\prime}})=g(v^{k}_{\ell}).$
We can then conclude that $\operatorname{BNN}(v^{k}_{\ell})=g(v^{k}_{\ell}).$
@@
$\operatorname{BNN}_{\sigma_{k}}(p)=\sum_{i=0}^{d}t^{k}_{i}\cdot g(v^{k}_{i}).$
If $p\in|\sigma^{i}|\cap|\sigma^{j}|$ for some $i,j\in\{0,1,\dots,n\}$, let $(t^{i}_{0},t^{i}_{1},\dots,t^{i}_{d})$ and $(t^{j}_{0},t^{j}_{1},\dots,t^{j}_{d})$ be the barycentric coordinates of $p$ with respect to $\sigma_{i}$ and $\sigma_{j}$ respectively. By the definition of simplicial complexes, there is an $\ell$-simplex $\sigma=\{v_{0},v_{1},\dots,v_{\ell}\}\subset\sigma_{i},\sigma_{j}$ with $\ell<d$ such that $p\in|\sigma|$ and satisfying that there are sets of indices $I,J\subset\{0,1,\dots,d\}$ such that $v_{r}=v^{i}_{r_{i}}=v^{j}_{r_{j}}$ and $t_{r}=t^{i}_{r_{i}}=t^{j}_{r_{j}}$ for some $r_{i}\in I$ and $r_{j}\in J$ and for all $r\in\{0,1,\dots,\ell\}$, being $(t_{0},t_{1},\dots,t_{\ell})$ the barycentric coordinates of $p$ with respect to $\sigma$. Putting all together,
$\operatorname{BNN}_{\sigma^{i}}(p)=\sum_{r_{i}\in I}t^{i}_{r_{i}}\cdot g(v^{i}_{r_{i}})=\sum_{r_{j}\in I}t^{j}_{r_{j}}\cdot g(v^{j}_{r_{j}})=\operatorname{BNN}_{\sigma^{j}}(p).$
We then conclude that, for any $k\in\{0,1,\dots,n\}$ with $p\in|\sigma_{k}|$, $\operatorname{BNN}(p)=\sum_{i=0}^{d}t^{k}_{i}\cdot g(v^{k}_{i})$, where $(t^{k}_{0},t^{k}_{1},\dots,t^{k}_{d})$ are the barycentric coordinates of $p$ with respect to $\sigma_{k}$.
If $p\not\in C$ then, $p\not\in|\sigma_{k}|$ for any given $k\in\{0,1,\dots,n\}$. Therefore, by Lemma 1, $\operatorname{BNN}_{\sigma^{k}}(p)=0$ and so $\operatorname{BNN}(p)=0$.
@@
where $m_{k}\in\mathbb{R}^{d}$ and $c_{k}\in\mathbb{R}$, concluding the proof. ∎
A key constraint in this setup is the following:
@@
Given a fixed number $n$, a BNN with $n$ base points can only approximate functions $f:[A,B]\to\mathbb{R}$ (sampled by a point cloud $X$) whose persistence barcode contains at most $n/2$ bars, or, in the noisy case, the $n/2$ most significant bars.
Still, BNNs are universal approximators.
Remark 6.
Any $\operatorname{CPLF}$ $g:[A,B]\to\mathbb{R}$ can be represented by a $\operatorname{BNN}$, with the key property that the $\operatorname{BNN}$ does not require an analytic definition. Instead, to specify the $\operatorname{CPLF}$ or the corresponding $\operatorname{BNN}$, it is sufficient to provide a finite set of transition (base) points $\{a_{i}\}\subset\mathbb{R}$, where the $\operatorname{CPLF}$ shifts between linear segments, together with the $\operatorname{CPLF}$ value at these points. This allows the $\operatorname{BNN}$ to capture the contribution of each segment relative to the input, thereby generating the correct output. For consistency with the terminology introduced in the Introduction (Section 1), in the following we shall refer to these transition points as base points.